OpenStudy (anonymous):

PLEASE HELP ME !!!!! Why is g(x)=x^3/sin(x) an odd function ?

OpenStudy (anonymous):

Well, honestly I'm confused. http://en.wikipedia.org/wiki/Even_and_odd_functions By the explanation in here, an odd function is a function so: \[ f(-x) = -f(x) \] for every X in the function's domain. let's try: \[ g(x) = \frac{x^3}{sin(x)}\\ g(1) = \frac{1^3}{sin(1)} = \frac{1}{sin(1)}\\ g(-1) = \frac{(-1)^3}{sin(-1)} = \frac{-1}{-sin(1)} = \frac{1}{sin(1)} \] And... \[ \frac{1}{sin(1)} \ne -\bigg[ \frac{1}{sin(1)} \bigg] \] So.. \[ g(-1) \ne -g(1) \] And therefore it can't be odd function by the definition... it's actually pretty easy to see that it is even function as \[ g(x) = g(-x) \] in here.. They also write there: "Examples of odd functions are x, x3, sin(x), sinh(x), and erf(x)." And "The quotient of two odd functions is an even function."

OpenStudy (anonymous):

but if you graph the eqution it looks like it is on the origin which is an odd function

OpenStudy (anonymous):

You mean it goes through (0,0)?

OpenStudy (anonymous):

I don't see it that way O.o... sin(0) should be... 0.... and we can't divide by 0. maybe it should be cos(x) instead?

OpenStudy (anonymous):

well if you do the odd function test g(-x)

OpenStudy (anonymous):

and the -g(x)

OpenStudy (anonymous):

like i did above? besides \( x \ne 0 \) and eh... http://www.wolframalpha.com/input/?i=x%5E3%2Fsinx They even say there its parity is 'even'.. idk what else you mean then

OpenStudy (anonymous):

Oh ! yeah thank you !

OpenStudy (anonymous):

I just wanted to see how to come about the answer algebraically

OpenStudy (anonymous):

so what would be f(x)= 0.001x^5+10x-57 ?

OpenStudy (anonymous):

Ok. \[ x^3 = x \cdot x \cdot x \\ (-x)^3 = (-x) \cdot (-x) \cdot (-x) = -(x \cdot x \cdot x) \\ \text{Therefore:} \\ -(x^3) = (-x)^3 \quad \quad \to \quad\quad-(x \cdot x \cdot x) = -(x \cdot x \cdot x) \] Now... it's known that, and easy to see also that... \[ sin(-x) = -sin(x) \] Therefore they both are odd functions... means if \(x < 0 \) they would produce negative value.. and negative/negative is positive... so \[ g(-x) = \frac{(-x)^3}{sin(-x)} = \frac{x^3}{sin(x)} = g(x) \\ \text{So:} \\ g(-x) = g(x) \] Means... g(x) is even.

OpenStudy (anonymous):

I see ! so would this equation be neither f(x)=0.001x^5+10x-57

OpenStudy (anonymous):

\[ f(x) = 0.001x^5+10x-57 \] well.. I guess it would be neither.. since \[ f(1) = 0.001+ 10 - 57 = 0.001 - 47 = -46.999 \\ f(-1) = -0.001 - 10 - 57 = -(0.001 + 67) = -67.001\\ -(-46.999) \ne -67.001\\ 46.999 \ne -67.001\\ -f(1) \ne f(-1) \] Besides... http://www.wolframalpha.com/input/?i=+0.001x%5E5%2B10x-57 No parity mentioned

OpenStudy (anonymous):

I should also add for even, very sorry \[ -46.999 \ne -67.001 \\ f(1) \ne f(-1) \] There for this is not even and not odd(above).

OpenStudy (anonymous):

Okay thank you so much !!!

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

Nope :) thanks I think I understand now

OpenStudy (anonymous):

Sure np, good luck then